Ticket Number: HEA-NOC/20060413-1 Ticket Status: UPDATE
Ticket Type: unscheduled Resolver:
Ticket Opened: 20060413 13:51 UTC+1 Problem Start: 20060413 13:21 UTC+1
Ticket Update: 20060517 17:00 UTC+1 Problem End:
Site/line: St. Angela's
There is an unscheduled outage to St. Angela's College, Sligo.
20060413 13:49 CMC Attempted to contact customer, will retry at 14:00 -
during STACS business hours.
20060413 14:26 CMC Contacted Customer who confirm there appear to be no
problems with equipment on-site. Contacted BT who have issues ticket
20060413 15:46 CMC Connectivity restored after link between BT equipment
and HEAnet router is re-initialised on-site.
20060414 14:54 OMG St. Angela's link has gone down again. Contacted Paul
Foley but couldn't reach him. Contacted BT again; they have issued ticket
#1-285444211. Rang Damien Kearns in St. Angela's who said that Paul is out
of the country and that there wouldn't be any ICT staff back in until
Tuesday at the earliest (which also means that there is no one to
facilitate access to St. Angelas should BT request it).
20060414 15:39 OMG Dylan from BT called back to say that this is a
recurring problem with this circuit and that the cables need to be
reseated on the customer equipment anytime any part of the circuit goes
down. BT have parked the ticket until Tuesday when someone is back in St.
Angelas to reseat the cables.
20060417 18:13 AH Connectivity returns at 17:55.
20060420 15:30 PC We have been seeing intermittent bursts of CRC and input
errors on the ATM switch in Galway, on incoming traffic from BT to our
Have informed BT and asked them to re-check their equipment to see if they
also see errors.
20060420 16:40 PC - BT have confirmed that they are seeing bit errors on
their radio equipment in St. Angelas and have suggested swapping out an
interface on the radio kit there. This is likely to require downtime.
Assuming no further deterioration BT hope to make the change on Monday. I
have emailed an update to the main contacts at St Angelas and have
arranged to speak to the BT enginner tomorrow to obtain further details of
the work, including the time period required.
20060421 09:05 PC - Connectivity to St Angelas was lost at 07:57 this
morning. Have informed BT and requested that they send out an engineer as
soon as possible. Haven't been able to reach Damien yet, but have left a
voice message suggesting reseating the cable to try to restore
connectivity until an engineer can get to site.
20060421 09:55 PC - St Angelas rebooted their router around 09:37, and
that seems to have restored connectivity. However we are still seeing
input errors on our galway atm switch due to traffic we receive from the
St Angelas router (via BT).
20060421 12:04 PC - BT will have an engineer on site at 4pm today and also
have an engineer tomorrow if that proves necessary. Paul Foley will be
available on site and I have passed on his contact details to BT.
20060422 21:21 SOB Transistions continued during the day continuing until
19:17 when it went down and has remained down since then.
20060424 09:07 PC - Link transitions have continued to take place over the
weekend. BT replaced the radio card at St Angelas, but were unable to
climb the radio mast as weather conditions made it unsafe to do so. A BT
Engineer will again be on site this morning to perform further checks.
20060424 15:21 PC - BT have notified ourselves and St Angelas that they
will be taking the radio link down for 1 hour while testing takes place.
They are curently waiting for winds to die down before beginning this work
on the mast.
20060424 16:40 PC - BT have brought the circuit down at 15:30 and are
continuing to work on it (no user traffic can be sent). Despite replacing
the interface (on Friday) and upgrading the radio software (this morning)
BT are still seeing 'severely errored seconds' and 'frame loss seconds'
errors on their radio equipment. So far they have been unable to ascend
the mast due to rain and have been checking the fibre instead.
20060424 18:38 PC - Having tested the fibre and isolated sections of the
network, BT believe that an intermediate radio mast may be causing the
issues. In addition to the two endpoints, two intermediate outdoor radio
units are used as direct line of sight is not possible end-to-end. Wet
weather has prevented the replacement of the faulty unit so far. I have
contacted Paul Foley with an update and have asked the BT engineer to do
so as soon as he is finished this evening. This is so that Paul can make
a call as to whether, if the replacement work is not possible this
evening, whether the circuit should be restored to its previous state (as
it was earlier today). Currently the link is down.
20060425 09:07 PC - BT commenced changing the Outdoor unit (ODU) around
08:15 this morning. The circuit is currently still down.
20060425 09:34 PC - The ODU RF interface has been replaced but BT
currently have the fibres attached into test equipment to verify the fault
has been rectified. This is expected to take an hour during which the
circuit will remain down.
20060425 10:52 PC - BT have run test across the replaced ODU RF interface
and they are no longer seeing errors on that span. They are now going
back to StAngelas to perform full end-to-end tests which they have
estimated to take up to 2 hours.
Andrew expects to arrive in 10 mins and I have left a message for Paul
Foley to this effect and will continue to try to contact Paul.
20060425 12:00 PC - After sucessfully testing the link to full capacity,
BT reconnected the link at 11:45. Our own tests have also run without
errors. The interface counters on the ST Angelas and Galway ATM switch
have been cleared and we will continue to monitor the link.
20060425 14:03 JH BT Ireland called to report the circuit has been clear
of errors for 2 hours.
20060503 12:40 PC - A small number of errors (1x input error & 5x CRC
errors) occured at around 4pm on 25th April, which was the day the link
was reestablished. No other errors have occured over the last eight days.
20060511 14:10 BB The circuit transitioned down/up at 14:09 -
investigating error counts and possible reasons
20060516 17:29 PC - The circuit transitioned down/up at 16:20, recovering
after 17 seconds. Unable to contact the customer at the moment to
determine if work was taking place, so have copied & then cleared router
20060517 12:51 PC - No work that could have affected the circuit was being
carried out in St Angelas yesterday. Have logged the call as a P2 with BT
Ireland, their ref: 1-311500827.
20060517 16:48 PC - BT identified "low receive errors" on the radio link
and have indicated these are likely to be caused by weather conditions,
such as rain. The radio team will continue monitoring the circuit over
the next few days.
This ticket can be monitored at http://www.hea.net/tickets/20060413-1